Ehrenberg and Smith Chapter 2 Problems 2 and 4

2. Suppose that the supply curve for schoolteachers is LS = 20,000 + 350W, and the demand curve for schoolteachers is LD = 100,000 - 150W, where L = the number of teachers and W = the daily wage.

a. Plot the supply and demand curves.

b. What are the equilibrium wage and employment levels in this market?

c. Now suppose that at any given wage, 20,000 more workers are willing to work as schoolteachers. Plot the new supply curve, and find the new wage and employment level. Why doesn't employment grow by 20,000?

4. Suppose the adult population of a city is 9,823,000 and there are 3,340,000 people who are not in the labor force and 6,094,000 who are employed.

a. Calculate the number of adults who are in the labor force and the number of adults who are unemployed.

b. Calculate the labor force participation rate and the unemployment rate.
Answer #1

2. a.

To plot the labor demand and supply curves, let us calculate demand and supply by giving random values to W.

If W = 0, then demand will be 100000 and supply will be 20000.

If W = 200, then demand will be 70000 and supply will be 90000. Plotting these values will give us the respective curves.

b.Solving the labor supply and demand curves we get

100000-150W = 20000+350W

100000-150W-20000-350W=0

80000-500W=0

W=160

L = 20000+350W

L = 20000+350*160

L = 20000+56000

L = 76000

c. To plot the new supply curve, let us calculate supply by giving random values to W as done in part a. The new supply curve will be 40000+350W

If W = 0, then supply will be 40000.

If W = 200, then supply will be 110000. Plotting these values will give us the respective curves. The new supply curve is the ns line in the below figure.

Solving the new labor supply and demand curves we get

100000-150W = 40000+350W

100000-150W-40000-350W=0

60000-500W=0

W=120

L = 40000+350W

L = 40000+350*120

L = 40000+42000

L = 82000

The employment does not grow by 20000 because, with increase in supply (shift in the supply curve) the wages fall.forcing the teachers to drop out of the labor market.

4. a. Number of adults in the labor force = total population - people not in labor force

= 9823000 - 3340000

= 6483000

Number of adults who are unemployed = total labor force - total employed

= 6483000 - 6094000

= 389000

b. Labor force participation rate = (labor force/population)*100

= (6483000/9823000)*100

= 66%

Unemployment rate = (total unemployed/total labor force)*100

= (389000/6483000)*100

= 6%

250 200 150 100 50 0 20000 40000 60000 80000 100000 120000

250 200 150 100 ns 50 0 20000 40000 60000 80000 100000 120000